Abbeville Fixture Company manufactures units in a small manufacturing facility. The units are made from brass. Manufacturing has 30 employees. Each employee presently provides 40 hours of labor per week. Information about a production week is as follows:

Standard wage per hour $10.8
Standard labor time per unit 10 min.
Standard number of lbs. of brass 1.8 lbs.
Standard price per lb. of brass $12.75
Actual price per lb. of brass $13
Actual lbs. of brass used during the week 15,574 lbs.
Number of units produced during the week 8,400
Actual wage per hour $11.12
Actual hours for the week (30 employees × 40 hours) 1,200 hrs.

Required:

a.  Determine the standard cost per unit for direct materials and direct labor. Round the cost per unit to two decimal places.

Direct materials standard cost per unit $
Direct labor standard cost per unit $
Total standard cost per unit $

b.  Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Materials Price Variance $
Direct Materials Quantity Variance $
Total Direct Materials Cost Variance $

c.  Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance $
Direct Labor Time Variance $
Total Direct Labor Cost Variance $
Answer #1

a. Direct materials standard cost per unit = 1.8*12.75 = $22.95 per unit

Direct labor standard cost per unit = 10/60*10.80 = $1.8 per unit

Total standard cost per unit= 22.95 + 1.80 = $24.75 per unit

b. Materials price variance = (Standard price - actual price) x Material purchased
= {$12.75 - $13} x 15574= $3894 Unfavorable

Materials quantity variance = (Standard quantity - actual quantity used) x standard price
= {(8400*1.8) - 15574} x $12.75 = $5789 Unfavorable

Total Direct Materials Cost Variance = 8400*22.95 - 15574*13 = $9682 Unfavorable

c. Labor rate variance = (Standard rate - actual rate) x Actual hours used
= {$10.8 - $11.12} x 1200 = $384 Unfavorable

Labor efficiency variance = (Standard hours - actual hours) x Standard rate
= {(8400*10/60) - 1200} x $10.80 = $2160 Favorable

Total Direct Labor Cost Variance =8400*1.80 - 1200*11.12 = $1776 Favorable